Suppose there are 6 white and 5 black marbles in a bag and you want to select 2 marbles. What is the probability that you will select a white marble then a black marble? You will replace the white marble. [give your answer to 4 d.p.]


1
Expert's answer
2021-09-30T14:18:16-0400

Total number of marbles = 6+5=11

The probability of selecting white marble is

"P(white) = \frac{6}{11}"

Since the marble is replaced, so total number of marbles will remain same.

So, the probability of selecting black marble is

"P(black) = \frac{5}{11}"

Therefore, the probability of selecting a white marble then a black marble with replacement is calculated as follows:

"P(white \; then \; black) = P(white) \times P(black) \\

= \frac{6}{11} \times \frac{5}{11} \\

= \frac{30}{121} =0.2479"